OpenStudy (rock_mit182):

A 0.40-kg object connected to a light spring with a force constant of 19.6 N/m oscillates on a frictionless horizontal surface. If the spring is compressed 4.0 cm and released from rest, determine (a) the maximum speed of the object, (b) the speed of the object when the spring is compressed 1.5 cm, and (c) the speed of the object when the spring is stretched 1.5 cm. (d) For what value of x does the speed equal one-half the maximum speed?

OpenStudy (anonymous):

since the spring is compressed by 4cm , amplitude is A = 4cm as the total energy is constant ,max. KE = max . PE (a) therefore (.5)*m*v^2 = (.5)*19.6*A^2 this will give max. speed (b) again use conservation of energy ,the total energy is (.5)*19.6*A^2 total energy = KE + PE at any instant (C) same as above ,ans will be same too (D) again by conservation of mechanical energy , KE at this instant is (.25) times of maxmimum kinetic energy which you already know

OpenStudy (rock_mit182):

thanks a lot bro!